OpenStudy (anonymous):

In a scale drawing of Samantha's apartment, the ratio is 1 in = 2 ft. The length of the living room on the drawing is 6 in and the width is 4 in. What is the perimeter of Samantha's living room?

OpenStudy (anonymous):

A 20 ft B 40 ft C 33 ft

jimthompson5910 (jim_thompson5910):

1/2 = 6/L 1*L = 6*2 L = 12 So her living room has an actual length of 12 ft

jimthompson5910 (jim_thompson5910):

1/2 = 4/W solve this for W and tell me what you get

OpenStudy (anonymous):

I don't know what to do

jimthompson5910 (jim_thompson5910):

cross multiply to get 1*W = 2*4 W = ???